OpenStudy (anonymous):

rlly need help with this one! 2x2 + 16x + 30 divided by 5x2 + 13x -6

OpenStudy (anonymous):

2x2 + 16x + 30 =2 (x^2 +8x + 15) = 2 (x +5) (x +3) 5x^2 + 13x -6 = ( x + 3) ( 5x-2) => 2 (x +5) (x +3) / ( x + 3) ( 5x-2) = 2 (x +5) / ( 5x-2)

OpenStudy (anonymous):

another form \[\frac{2x+10}{5x-2}\]

OpenStudy (anonymous):

Thanks for giving out the medal!, Marina :)

OpenStudy (anonymous):

you are doing good job and deserve reward!

OpenStudy (anonymous):

I have no clue about who give whom, as I'm brand new member, for 2 days.

OpenStudy (anonymous):

You got 179 medals for 2 days???

OpenStudy (anonymous):

I guess there are so many newbie learning Math, so I keep solving for them :)

OpenStudy (anonymous):

you really dedicated one :-)))